According to a recent​ study, 68% of adults say they play games on their smartphone.

Suppose we randomly select 200 adults and find that​ 80% of the sample plays games on

their smartphone. Which of the following correctly describes the distribution for the

proportion of adults in the sample who play games on their​ smartphone?

A. ​pˆ ~​ AN(0.68, 0.0283)

B. pˆ ~​ AN(0.68, 0.0330)

C. X​ ~ AN(160,​ 5.6569)

D. ​pˆ ~​ AN(0.80, 0.0283)
Answer #1

Answer: B. pˆ ~​ AN(0.68, 0.0330)

Solution:

We are given

n = 200

p = 0.68

q = 1 - 0.68 = 0.32

Estimate for mean for sampling distribution of sample proportion µp̂ = p = 0.68

µp̂ = 0.68

Estimate for standard deviation for sampling distribution of sample proportion σp̂ is given as below:

σp̂ = sqrt(pq/n) = sqrt(0.68*0.32/200) = 0.032985

σp̂ = 0.033

Answer:

B. pˆ ~​ AN(0.68, 0.0330)